# Multilingual Performance Biases of Large Language Models in Education

### v1

@@ Abstract
Large language models (LLMs) are increasingly being adopted in educational settings. These applications expand beyond English, though current LLMs remain primarily English-centric. In this work, we ascertain if their use in education settings in non-English languages is warranted. We evaluated the performance of popular LLMs on four educational tasks: identifying student misconceptions, providing targeted feedback, interactive tutoring, and grading translations in six languages (Hindi, Arabic, Farsi, Telugu, Ukrainian, Czech) in addition to English. We find that the performance on these tasks somewhat corresponds to the amount of language represented in training data, with lower-resource languages having poorer task performance. Although the models perform reasonably well in most languages, the frequent performance drop from English is significant. Thus, we recommend that practitioners first verify that the LLM works well in the target language for their educational task before deployment.
@@
Multilingual Performance Biases of Large Language Models in Education
Vansh ${}^{\bigstar}{}^{\dagger}$   Sankalan Pal ${}^{\bigstar}{}^{\dagger}$ Vilém Zouhar   Donya Rooein   Mrinmaya Sachan { guptav, spalchowd, vzouhar, msachan}@ethz.ch   donya.rooein@unibocconi.it ETH Zurich  Bocconi University
footnotetext: Co-first authors sorted alphabetically. 00footnotetext: We release the collected dataset and code publicly at github.com/eth-lre/multilingual-educational-llm-bias. The dataset comprises 91,000 automatically evaluated model outputs across six languages, four tasks, and five models.
@@ 1 Introduction
Although most of these LLMs are trained on multilingual corpora (OpenAI, 2019; Nvidia, 2022; Peng et al., 2023; Gu and Dao, 2023), they are still overwhelmingly English-centric (Argoub, 2022; Ruder et al., 2022, Table 1). Inadequate adaptation to local languages in an educational setting risks diminishing their utility and exacerbating existing inequalities by privileging dominant languages and cultures. The question of multilingualism arises in every domain where LLMs are applied Lai et al. (2023); Ahuja et al. (2023, 2024). Still, it is imperative in the field of education, which has seen wide use of LLMs despite the high stakes (Alhafni et al., 2024; Raheja et al., 2023; Naismith et al., 2023). Without rigorous evaluation tailored to educational tasks across languages, deploying LLMs in classrooms may introduce new forms of harm, including misinformation, misalignment with curricula, or culturally inappropriate content (Almasoud et al., 2025).
In this work, we present an empirical investigation of the capabilities of frontier LLMs on educational tasks across several languages. We identify four education-related tasks (identifying student misconceptions, providing targeted feedback, interactive tutoring, and translation grading) with well-defined language-agnostic metrics. We then evaluate several frontier LLMs (Claude, Gemini, GPT4o, Llama, and Mistral) on these tasks in six languages (Hindi, Arabic, Farsi, Telugu, Ukrainian, and Czech) in addition to English.
@@
Farsi Indo-Iranian Abjad 1034K 0.74% 127M
@@ 2 Methods
An important aspect of teaching is fixing student misconceptions, which first requires identifying the student misconception (Liu et al., 2023). We build this task on the EEDI Math Questions Dataset, which contains thousands of multiple-choice questions with four answer choices. For many of the wrong choices, we have expert-annotated misconceptions that could lead to a student picking the said choice. We leverage these to build our task. The LLM is given a multiple-choice question, the student’s (incorrect) answer, and four possible misconceptions. The candidate misconceptions include the true misconception identified by experts and three distractors chosen at random from the other misconceptions present in the dataset. The LLM must pick the correct misconception from these four options (see Figure 5 for an example). We evaluate the LLM performance by reporting accuracy in predicting the student misconception. Since the model must pick one of four options, a random baseline has an accuracy of 25%.
@@
A common field of education that has seen an increase in the use of LLMs is Language learning Klimova et al. (2024); Zhu et al. (2024). A representative task from this field is to assign a grade to a translation provided by a student. While we lack proper datasets across languages with translations and their appropriate grades, we can approximate this task by the fact that the machine translation of a sentence should receive a higher grade than the exact translation with one word replaced by a random word.
@@
We choose six languages for experiments: Hindi, Telugu, Farsi, Ukrainian, and Czech, in addition to English for comparison. This language selection reflects diverse linguistic properties, varying levels of representation in training data, and different language families (Foundation, 2024), see Table 1. Hindi (Indo-Aryan) and Telugu (Dravidian) represent major languages from the Indian subcontinent that use the Brahmic script and are under-represented in both CommonCrawl and Wikipedia. Farsi and Arabic offer insights into LLM performance on a right-to-left Abjad script, whereas Ukrainian and Czech allow us to study generalisation in medium resource morphologically rich languages, using the Cyrillic and Latin scripts, respectively.
@@
6.079
A: Confuses powers and multiples,
B (correct): When multiplying a surd by an integer to write as one surd, squares the number under the surd rather than the number in front,
C: When ordering integers, orders from the digits at the end instead of the digits at the start,
D: Does not understand that when multiplying both sides of an equation by an amount every term must be multiplied by the same amount
@@
Student answer: $4$
@@
A (student answer): I think you have made an arithmetic error when halving £1.50. Use short division to divide by two,
@@
D: I think you have found the cost for one pencil. The question asks for the cost of 3 pencils.
@@
C (correct answer): Almost there! Take care to notice how many hours and minutes you’re adding here. Is your answer 2 hours and 52 minutes later than 8.50pm?,
D: Adding 2 hours to 8.50pm gives 10.50pm. Adding 10 minutes on takes us to 11.00pm, and adding the remaining 42 minutes gives 11.42pm.
@@
It is a kind of tomato. Refer to caption Refer to caption Hindi
@@
To assess the typological diversity of our selected languages, we used the URIEL typological database Littell et al. (2017) with lang2vec, which provides dense vector representations of languages based on a range of typological, phylogenetic, and geographical features. As recommended by the package, we extracted syntax features with k-NN predictions for the missing values for a set of 40 languages, constructed as the union of our core experimental languages and the most widely spoken languages worldwide according to Ethnologue Eberhard et al. (2025). We projected each language feature vector into two dimensions using Multidimensional Scaling, producing a 2D language similarity plot. This allows us to visualise (see Figure 1) the relative syntactic diversity of our selected languages and confirm that they span a broad typological space. The visualisation demonstrates that our language selection (highlighted) is well distributed across the typological landscape.
@@
Models and prompts.
We evaluate six state-of-the-art LLMs praised for their multilingual capabilities: GPT-4o OpenAI (2023), Gemini 2.0 Flash Team (2024), Claude 3.7 Sonnet Anthropic (2024), Llama 3.1 405B Grattafiori et al. (2024), Mistral Large 2407 AI (2024); Jiang et al. (2023), and Command-A Cohere et al. (2025). We leave all sampling parameters to their defaults. For prompts, we use a simple chain of thought prompting method, where the model is first asked to explain why it would pick a certain answer, and then asked to choose it in a separate prompt. Based on literature Mondshine et al. (2024); Huang et al. (2023), it is unclear whether or not it is beneficial to translate the prompt itself to the target language or keep it in English, so we try both options.111 A weaker model roleplays the student model used in the tutoring task to be consistent with the original work. We only use the original prompts because it does not work well with non-English prompts. , [^2] We machine-translate the prompts and manually verify (with L1/L2 language knowledge) the translation adequacy.
English prompt Translated prompt
Farsi $\cdot$ 94.8% $\cdot$ 93.3% $\cdot$ 93.0% $\cdot$ 87.5% $\cdot$ 92.7% $\cdot$ 93.1% $\cdot$ 95.1% $\cdot$ 94.4% $\bigstar$68.0$\cdot$ 88.3% $\bigstar$66.9$\cdot$ 93.6%
Telugu $\cdot$ 95.2% $\cdot$ 92.2% $\cdot$ 89.9% $\cdot$ 86.9% $\cdot$ 89.7% $\star$ 85.5% $\cdot$ 94.2% $\cdot$ 90.8% $\bigstar$68.6$\star$ 83.6% $\bigstar$35.5$\bigstar$77.9%
Ukrainian $\cdot$ 95.7% 94.9% $\cdot$ 92.9% 93.3% 94.4% 94.9% $\cdot$ 95.6% $\cdot$ 94.3% $\bigstar$56.6$\cdot$ 90.4% 94.2% 93.9%
Czech 96.9% 95.1% 94.5% 92.3% 94.5% 94.1% 96.6% 95.8% $\bigstar$70.2$\star$ 81.6% $\bigstar$41.094.5%
Table 3: Results (accuracy) for the misconception identification task. We mark results significantly lower (at least 10%$\bigstar$, at least 5%$\star$, otherwise $\cdot$) than English with a one-sided 95% confidence t-test.
English prompt Translated prompt
Hindi $\cdot$ 48.7% 35.6% $\cdot$ 13.0% $\cdot$ 43.6% $\cdot$ 40.5% $\cdot$ 31.6% $\bigstar$32.1$\bigstar$13.4$\star$ 6.2% $\cdot$ 44.3% $\bigstar$18.6$\bigstar$18.8%
Arabic $\cdot$ 49.6% $\star$ 28.7% $\cdot$ 13.9% $\cdot$ 45.3% $\star$ 38.8% $\cdot$ 33.3% $\cdot$ 48.8% $\bigstar$10.716.3% 48.1% $\bigstar$27.8$\star$ 28.9%
Farsi 50.2% $\star$ 27.9% $\cdot$ 11.3% $\cdot$ 44.9% $\cdot$ 41.3% $\star$ 30.9% $\cdot$ 45.9% $\cdot$ 31.6% 16.3% $\cdot$ 44.0% $\bigstar$33.5$\cdot$ 35.5%
Telugu $\cdot$ 45.2% $\star$ 27.6% $\cdot$ 10.4% $\cdot$ 43.4% $\bigstar$34.0$\star$ 26.3% $\bigstar$13.9$\bigstar$12.7$\star$ 6.1% $\star$ 37.7% $\bigstar$15.5$\bigstar$9.5%
Ukrainian 50.3% $\cdot$ 33.2% $\cdot$ 13.0% $\cdot$ 44.8% $\cdot$ 41.3% $\cdot$ 32.2% $\bigstar$35.9$\bigstar$19.6$\star$ 8.1% 52.8% $\bigstar$31.0$\star$ 27.2%
Table 4: Results (accuracy) for the feedback selection task. We mark results significantly lower (at least 10%$\bigstar$, at least 5%$\star$, otherwise $\cdot$) than English with a one-sided 95% confidence t-test.
Harmonic mean Success/1-Telling
Telugu $\bigstar$50.1$\bigstar$39.527.7% $\cdot$ 58.9% $\bigstar$2.9$\bigstar$40.777.5/40.5% 77.5/51.0% 85.5/69.0% 61.0/4.0% 59.0/57.5% 63.5/33.5%
Ukrainian 91.2% 91.5% 23.5% $\cdot$ 81.2% 71.5% 90.9% 93.0/3.5% 92.0/1.0% 91.5/78.0% 84.0/5.5% 71.5/0.0% 93.5/5.0%
Czech $\bigstar$43.8$\bigstar$44.117.2% 70.2% $\bigstar$2.9$\bigstar$21.565.5/32.5% 73.5/42.0% 90.0/80.5% 71.5/2.5% 52.5/51.0% 77.0/64.5%
Table 5: Results (harmonic mean, success, and telling) for the tutoring task. We mark results significantly lower (at least 10%$\bigstar$, at least 5%$\star$, otherwise $\cdot$) than English with a one-sided 95% confidence t-test when occurring in both success and telling. Telling is flipped such that higher is better.
English prompt Translated prompt
Ukrainian 98.0% 97.3% 96.9% 96.5% 97.3% 98.3% 98.1% 97.9% 85.3% 97.7% 98.4% 98.2%
Table 6: Results (accuracy) for the translation grading task.
@@ 3 Results
In this section, we describe the results of five popular large language models on the four tasks described in Section 2. The main results are shown in Tables 6, 6, 6 and 6.
@@
The gap between English and other languages is large in general. On average across all tasks (excluding translation) and models, English has 70.9%, in contrast to 63.1% (Hindi), 55.3% (Czech), 67.8% (Ukrainian), 49.7% (Telugu), 66.2% (Farsi) and 67.4%(Arabic). This in itself does not make it clear if the loss is due to the LLMs being weak or the translation quality being poor. The poor performance on Telugu is largely driven by Command-A and Mistral. The former is unsurprising as Telugu is the only language in our list that is not officially supported by it Cohere et al. (2025). On the other hand, Mistral lists only 12 supported languages of which we test only Hindi and Arabic. Telugu also has the lowest representation in CommonCrawl and Wikipedia, so the result is expected. Manual analysis of the low tutoring performance for Czech reveals that the interactions switch between various language formality styles, to the point that it becomes distracting. Additionally, the language used in Czech classrooms is particular and likely not represented on the internet.
@@
Mistral is the most inconsistent across non-English languages (average deviation [^3] We calculate the standard deviation across the six languages for each task and then calculate the mean.=0.204). For example, it completely fails the tutoring task for both Czech and Telugu, despite performing reasonably on other languages in the same task. Command-A is not much better (average deviation3=0.179). Further, Command-A continues to be the second-worst on the list should we exclude Telugu from our analysis. On the other hand, Gemini is the most consistent (average deviation3=0.075) and also has the second-best performance (average score 74.3%). GPT4o, is the best performing model (average 76.9%) while Claude performs the worst (average 50.0%) mostly due to Feedback and Tutoring tasks.
@@
The worst performance is observed in the Feedback task despite the similarity to the Misconception identification task. While Claude is still the standout worst performer with a worse-than-random performance, all models struggle. Further analysis in Table 11 shows that all models tended to default the feedback corresponding to the correct answer, with the models’ chain of thoughts being “regardless of the student’s mistake, this is the feedback that gives the student the most information about the correct answer. Most models perform well in the Translation evaluation task, with the accuracy being even higher than human annotators, who were presented with attention checks with similar perturbations (Kocmi et al., 2024; Zouhar et al., 2025). They also do well in the Misconceptions task, with most percentage scores (at least in the English prompt setting) being in the 90s. The tutoring task seems to have the most inconsistent performance across models and languages. In general, all models struggle in Czech and Telugu, while Claude struggles in all languages. Avoiding telling seems to be the more challenging part of the problem for all the models, although success rates are not very consistent either.
@@
Except for the Tutoring task, using English prompts yields better performance than using translated prompts (averages 71.6% and 66.2%). The exceptions to these are GPT, Llama, Gemini, and Mistral in the translation task, and Gemini in the Feedback task, though in most cases, the difference is not very large. Note that some of the poor performance could be attributed to the prompts being translated and checked for correctness rather than being written in the target language directly, which could introduce some translationese. Regardless, we believe it is best to keep prompts in English. As a further note for English-speaking developers designing multilingual applications, keeping prompts in English ensures that the chains-of-thought remain English, making it easier to run sanity checks.

### v2

@@ Abstract
Large language models (LLMs) are increasingly being adopted in educational settings. These applications expand beyond English, though current LLMs remain primarily English-centric. In this work, we ascertain if their use in education settings in non-English languages is warranted. We evaluated the performance of popular LLMs on four educational tasks: identifying student misconceptions, providing targeted feedback, interactive tutoring, and grading translations in eight languages (Mandarin, Hindi, Arabic, German, Farsi, Telugu, Ukrainian, Czech) in addition to English. We find that the performance on these tasks somewhat corresponds to the amount of language represented in training data, with lower-resource languages having poorer task performance. However, at least some models are able to more or less maintain their levels of performance across all languages. Thus, we recommend that practitioners first verify that the LLM works well in the target language for their educational task before deployment.
@@
Are Large Language Models for Education Reliable for All Languages?
[…]
00footnotetext: We release the collected dataset and code at github.com/ eth-lre/multilingual-educational-llm-bias. The dataset comprises 313,500 automatically evaluated model outputs across seven languages, four tasks, and six models.
@@ 1 Introduction
Although most of these LLMs are trained on multilingual corpora (OpenAI, 2019; Nvidia, 2022; Peng et al., 2023; Gu and Dao, 2023), they are still overwhelmingly English-centric (Argoub, 2022; Ruder et al., 2022, Table˜1). Inadequate adaptation to local languages in an educational setting risks diminishing their utility and exacerbating existing inequalities by privileging dominant languages and cultures. The question of multilingualism arises in every domain where LLMs are applied Lai et al. (2023); Ahuja et al. (2023, 2024). However, it is especially important in the field of education, which has seen wide use of LLMs despite the high stakes (Alhafni et al., 2024; Raheja et al., 2023; Naismith et al., 2023). Without rigorous evaluation tailored to educational tasks across languages, deploying LLMs in classrooms may introduce new forms of harm, including misinformation, misalignment with curricula, or culturally inappropriate content (Almasoud et al., 2025).
In this work, we present an empirical investigation of the capabilities of frontier LLMs on educational tasks across several languages. We identify four education-related tasks (identifying student misconceptions, providing targeted feedback, interactive tutoring, and translation grading) with well-defined language-agnostic metrics. We then evaluate several frontier LLMs (Claude, Gemini, GPT4o, Llama, and Mistral) on these tasks in eight languages (Mandarin, Hindi, Arabic, German, Farsi, Telugu, Ukrainian, and Czech) in addition to English.
@@
Mandarin Sino-tibetan Hanzi [^1] Alternately referred to as Kanji, Hanja or Hantu 1480K 5.8% 1184M
@@
German Germanic Latin 3021K 5.5% 411M
Farsi Indo-Iranian Abjad 1034K 0.74% 134M
@@ 2 Methods
An important aspect of teaching is fixing student misconceptions, which first requires identifying the student misconception (Liu et al., 2023). We build this task on the EEDI Math Questions Dataset, which contains thousands of multiple-choice questions with four answer choices. For many of the wrong choices, we have expert-annotated misconceptions that could lead to a student picking the said choice. We leverage these to build our task. The LLM is given a multiple-choice question, the student’s (incorrect) answer, and four possible misconceptions. The candidate misconceptions include the true misconception identified by experts and three distractors chosen at random from the other misconceptions present in the dataset. The LLM must pick the correct misconception from these four options (see Figure˜5 for an example). We evaluate the LLM performance by reporting accuracy in predicting the student misconception. Since the model must pick one of four options, a random baseline has an accuracy of 25%.
@@
Language Questions Misconception Feedback Translation
Mandarin 0.593 0.607 0.666 0.781
Hindi 0.455 0.546 0.593 0.831
Arabic 0.596 0.659 0.605 0.793
German 0.623 0.642 0.697 0.792
Farsi 0.574 0.644 0.708 0.840
Telugu 0.518 0.569 0.578 0.639
Ukrainian 0.607 0.642 0.682 0.821
Czech 0.611 0.626 0.663 0.805
[…]
@@
We choose eight languages for experiments: Mandarin, Hindi, Arabic, German, Farsi, Telugu, Ukranian, Czech in addition to English for comparison. This language selection reflects diverse linguistic properties, varying levels of representation in training data, and different language families (Foundation, 2024), see Table˜1. Hindi (Indo-Aryan) and Telugu (Dravidian) represent major languages from the Indian subcontinent that use the Brahmic script and are under-represented in both CommonCrawl and Wikipedia. German and Mandarin, on the other hand, are examples of languages well represented in both CommonCrawl and Wikipedia Farsi and Arabic offer insights into LLM performance on a right-to-left Abjad script, whereas Ukrainian and Czech allow us to study generalisation in medium resource morphologically rich languages, using the Cyrillic and Latin scripts, respectively.
@@
5.0001
Right Answer:5.2
A: Believes the mean is total frequency divided by something,
B (correct): Thinks the more digits a number has the greater it is, regardless of place value,
C: Believes parallel lines have gradients that multiply to give -1,
D: When multiplying by a multiple of 10, gives an answer 10 times bigger than it should be
@@
Student answer: 44
Right Answer: 8
@@
A: I think you have made an arithmetic error when halving £1.50. Use short division to divide by two,
@@
D (student answer): I think you have found the cost for one pencil. The question asks for the cost of 3 pencils.
@@
C: Almost there! Take care to notice how many hours and minutes you’re adding here. Is your answer 2 hours and 52 minutes later than 8.50pm?,
D (correct answer): Adding 2 hours to 8.50pm gives 10.50pm. Adding 10 minutes on takes us to 11.00pm, and adding the remaining 42 minutes gives 11.42pm.
@@
It is a kind of tomato. Refer to caption Refer to caption Mandarin
Refer to caption Refer to caption Hindi
@@
Es ist eine Art Tomate Katze ist eine Art Tomate German
@@
To assess the typological diversity of our selected languages, we used the URIEL typological database Littell et al. (2017) with lang2vec, which provides dense vector representations of languages based on a range of typological, phylogenetic, and geographical features. As recommended by the package, we extracted syntax features with k-NN predictions for the missing values for a set of 40 languages, constructed as the union of our core experimental languages and the most widely spoken languages worldwide according to Ethnologue Eberhard et al. (2025). We projected each language feature vector into two dimensions using Multidimensional Scaling, producing a 2D language similarity plot. This allows us to visualise (see Figure˜1) the relative syntactic diversity of our selected languages and confirm that they span a broad typological space. The visualisation demonstrates that our language selection (highlighted) is well distributed across the typological landscape.
@@
Models and prompts.
We evaluate six state-of-the-art LLMs praised for their multilingual capabilities: GPT-4o OpenAI (2023), Gemini 2.0 Flash Team (2024), Claude 3.7 Sonnet Anthropic (2024), Llama 3.1 405B Grattafiori et al. (2024), Mistral Large 2407 AI (2024); Jiang et al. (2023), and Command-A Cohere et al. (2025). We leave all sampling parameters to their defaults. For prompts, we use a simple chain of thought prompting method, where the model is first asked to explain why it would pick a certain answer, and then asked to choose it in a separate prompt. Based on literature Mondshine et al. (2024); Huang et al. (2023), it is unclear whether or not it is beneficial to translate the prompt itself to the target language or keep it in English, so we try both options.222 A weaker model roleplays the student model used in the tutoring task to be consistent with the original work. We only use the original prompts because it does not work well with non-English prompts. , [^3] We machine-translate the prompts and manually verify (with L1/L2 language knowledge) the translation adequacy.
English prompt Translated prompt
Mandarin $\cdot$ 95.8% 95.2% $\cdot$ 92.5% $\cdot$ 92.1% $\cdot$ 92.8% $\cdot$ 92.9% 96.5% 95.0% $\cdot$ 91.7% 93.8% $\cdot$ 92.9% 94.1%
German $\cdot$ 96.0% 96.2% 94.6% $\star$ 84.6% 95.1% 95.2% $\cdot$ 95.9% 96.6% 94.0% $\bigstar$74.0% 94.9% 95.2%
Farsi $\cdot$ 94.8% $\cdot$ 93.3% $\cdot$ 93.0% $\cdot$ 87.5% $\cdot$ 92.7% $\cdot$ 93.1% $\cdot$ 95.1% $\cdot$ 94.4% $\bigstar$68.0% $\cdot$ 88.3% $\bigstar$66.9% $\cdot$ 93.6%
Telugu $\cdot$ 95.2% $\cdot$ 92.2% $\cdot$ 89.9% $\cdot$ 86.9% $\cdot$ 89.7% $\star$ 85.5% $\cdot$ 94.2% $\cdot$ 90.8% $\bigstar$68.6% $\star$ 83.6% $\bigstar$35.5% $\bigstar$77.9%
Ukranian $\cdot$ 95.7% 94.9% $\cdot$ 92.9% 93.3% 94.4% 94.9% $\cdot$ 95.6% $\cdot$ 94.3% $\bigstar$56.6% $\cdot$ 90.4% 94.2% 93.9%
Czech 96.9% 95.1% 94.5% 92.3% 94.5% 94.1% 96.6% 95.8% $\bigstar$70.2% $\star$ 81.6% $\bigstar$41.0% 94.5%
[…]
English prompt Translated prompt
Mandarin $\cdot$ 49.6% $\star$ 29.7% $\cdot$ 12.3% $\cdot$ 43.0% $\cdot$ 40.1% $\cdot$ 31.8% $\star$ 41.1% $\bigstar$19.2% $\star$ 5.8% $\bigstar$30.3% $\bigstar$30.3% $\star$ 27.8%
Hindi $\cdot$ 48.7% 35.6% $\cdot$ 13.0% $\cdot$ 43.6% $\cdot$ 40.5% $\cdot$ 31.6% $\bigstar$32.1% $\bigstar$13.4% $\star$ 6.2% $\cdot$ 44.3% $\bigstar$18.6% $\bigstar$18.8%
Arabic $\cdot$ 49.6% $\star$ 28.7% $\cdot$ 13.9% $\cdot$ 45.3% $\star$ 38.8% $\cdot$ 33.3% $\cdot$ 48.8% $\bigstar$10.7% 16.3% 48.1% $\bigstar$27.8% $\star$ 28.9%
German 52.5% $\cdot$ 32.1% 15.0% $\cdot$ 46.4% $\cdot$ 42.4% $\cdot$ 32.8% 50.6% $\cdot$ 30.8% 15.6% $\cdot$ 44.4% $\star$ 39.4% 37.6%
Farsi 50.2% $\star$ 27.9% $\cdot$ 11.3% $\cdot$ 44.9% $\cdot$ 41.3% $\star$ 30.9% $\cdot$ 45.9% $\cdot$ 31.6% 16.3% $\cdot$ 44.0% $\bigstar$33.5% $\cdot$ 35.5%
Telugu $\cdot$ 45.2% $\star$ 27.6% $\cdot$ 10.4% $\cdot$ 43.4% $\bigstar$34.0% $\star$ 26.3% $\bigstar$13.9% $\bigstar$12.7% $\star$ 6.1% $\star$ 37.7% $\bigstar$15.5% $\bigstar$9.5%
Ukranian 50.3% $\cdot$ 33.2% $\cdot$ 13.0% $\cdot$ 44.8% $\cdot$ 41.3% $\cdot$ 32.2% $\bigstar$35.9% $\bigstar$19.6% $\star$ 8.1% 52.8% $\bigstar$31.0% $\star$ 27.2%
[…]
Harmonic mean Success/1-Telling
Mandarin 89.8% 89.0% 26.4% 79.7% 79.7% 88.2% 94.0/8.0% 90.5/3.0% 90.0/74.5% 80.5/1.5% 80.0/0.5% 93.0/9.0%
German 90.7% 91.2% 23.4% 84.2% 77.2% $\cdot$ 86.3% 92.5/3.5% 92.0/1.5% 88.0/74.5% 85.0/1.5% 77.5/0.5% 90.5/8.0%
Telugu $\bigstar$50.1% $\bigstar$39.5% 27.7% $\cdot$ 58.9% $\bigstar$2.9% $\bigstar$40.7% 77.5/40.5% 77.5/51.0% 85.5/69.0% 61.0/4.0% 59.0/57.5% 63.5/33.5%
Ukranian 91.2% 91.5% 23.5% $\cdot$ 81.2% 71.5% 90.9% 93.0/3.5% 92.0/1.0% 91.5/78.0% 84.0/5.5% 71.5/0.0% 93.5/5.0%
Czech $\bigstar$43.8% $\bigstar$44.1% 17.2% 70.2% $\bigstar$2.9% $\bigstar$21.5% 65.5/32.5% 73.5/42.0% 90.0/80.5% 71.5/2.5% 52.5/51.0% 77.0/64.5%
[…]
English prompt Translated prompt
Mandarin 100.0% 99.3% 98.9% 99.9% 99.5% 99.9% 99.9% 99.4% 24.8% 99.6% 99.4% 99.9%
German 98.2% 97.9% 97.9% 98.2% 98.2% 98.2% 98.5% 98.3% 29.9% 98.0% 98.3% 97.8%
Ukranian 98.0% 97.3% 96.9% 96.5% 97.3% 98.3% 98.1% 97.9% 85.3% 97.7% 98.4% 98.2%
Table 6: Results (accuracy) for the translation grading task.
@@ 3 Results
In this section, we describe the results of five popular large language models on the four tasks described in Section˜2. The main results are shown in Tables˜6, 6, 6 and 6.
@@
The gap between English and other languages is large in general. On average [^4] Averaging here is done to give a general idea, but we must note that the scores are not equivalent. We use Accuracy for tasks 1 and 2 but Tutoring Score for Task 3 across all tasks (excluding translation) and models, English has 70.9%, in contrast to 63.1% (Hindi), 55.3% (Czech), 67.8% (Ukrainian), 49.7% (Telugu), 66.2% (Farsi), 66.8% (German), 64.6% (Mandarin) and 67.4% (Arabic). This in itself does not make it clear if the loss is due to the LLMs being weak or the translation quality being poor. The poor performance on Telugu is largely driven by Command-A and Mistral. The former is unsurprising as Telugu is the only language in our list that is not officially supported by it Cohere et al. (2025). On the other hand, Mistral lists only 12 supported languages of which we test only Hindi, Arabic, German and Chinese. Telugu also has the lowest representation in CommonCrawl and Wikipedia, so the result is expected. Manual analysis of the low tutoring performance for Czech reveals that the interactions switch between various language formality styles, to the point that it becomes distracting. Additionally, the language used in Czech classrooms is particular and likely not represented on the internet.
@@
Mistral is the most inconsistent across non-English languages (average deviation [^5] We calculate the standard deviation across the six languages for each task and then calculate the mean.=0.186). For example, it completely fails the tutoring task for both Czech and Telugu, despite performing reasonably on other languages in the same task. Command-A is not much better (average deviation5=0.161). On the other hand, Gemini is the most consistent (average deviation5=0.078) and also has the second-best performance (average score 75.0%). GPT4o, is the best performing model (average 78.6%) while Claude performs the worst (average 49.3%) mostly due to Feedback and Tutoring tasks.
@@
The worst performance is observed in the Feedback task despite the similarity to the Misconception identification task. While Claude is still the standout worst performer with a worse-than-random performance, all models struggle. Further analysis in Table˜12 shows that all models tended to default the feedback corresponding to the correct answer, with the models’ chain of thoughts being “regardless of the student’s mistake, this is the feedback that gives the student the most information about the correct answer.” Most models perform well in the Translation evaluation task, with the accuracy being even higher than human annotators, who were presented with attention checks with similar perturbations (Kocmi et al., 2024; Zouhar et al., 2025). They also do well in the Misconceptions task, with most percentage scores (at least in the English prompt setting) being in the 90s. The tutoring task seems to have the most inconsistent performance across models and languages. In general, all models struggle in Czech and Telugu, while Claude struggles in all languages. Avoiding telling seems to be the more challenging part of the problem for all the models, although success rates are not very consistent either.
@@
Excluding for the Tutoring task (which did not use native prompts), using English prompts yields better performance than using translated prompts (averages 72.7% and 67.2%). The exceptions to these are GPT, Llama, Gemini, and Mistral in the translation task though in most cases, the difference is not very large. Note that some of the poor performance could be attributed to the prompts being translated and checked for correctness rather than being written in the target language directly, which could introduce some translationese. Regardless, we believe it is best to keep prompts in English. As a further note for English-speaking developers designing multilingual applications, keeping prompts in English ensures that the chains-of-thought remain English, making it easier to run sanity checks.